A side of a square is 10 meters longer than the side of an equilateral triangle. The perimeter of the square is 3 times the perimeter of the triangle. Find the length of each side of the triangle.
Answer by mananth(16946)

square side = 10 meters
perimeter of square = 4*L = 40 meters
...
let side of triangle be x
perimeter of triangle = 3x
...
40 = 3x *3
9x =40
x= 40/9
x= 4.44 meters length of side of triangle
..
m.ananth@hotmail.ca